# A complete classification of the zero-dimensional homogeneous spaces under determinacy

### v2

@@ 3.1. The basics of Wadge theory
We begin with the most fundamental definitions and results of Wadge theory. For the proofs, see [CMM2, Section 4]. Assume that a set $Z$ and $\mathbf{\Gamma}\subseteq\mathcal{P}(Z)$ are given. We will use the notation $\widecheck{\mathbf{\Gamma}}=\{Z\setminus A:A\in\mathbf{\Gamma}\}$ for the dual of $\mathbf{\Gamma}$, and say that $\mathbf{\Gamma}$ is selfdual if $\mathbf{\Gamma}=\widecheck{\mathbf{\Gamma}}$. Also set $\Delta(\mathbf{\Gamma})=\mathbf{\Gamma}\cap\widecheck{\mathbf{\Gamma}}$.
@@
Let $Z$ be a space, and let $A,B\subseteq Z$. We will write $A\leq B$ if there exists a continuous function $f:Z\longrightarrow Z$ such that $A=f^{-1}[B]$. In this case, we will say that $A$ is Wadge-reducible to $B$, and that $f$ witnesses the reduction.111 Wadge-reduction is usually denoted by $\leq_{\mathsf{W}}$, which allows to distinguish it from other types of reduction (such as Lipschitz-reduction). Since we will not consider any other type of reduction in this article, we decided to simplify the notation. We will write $A<B$ if $A\leq B$ and $B\not\leq A$. We will say that $A$ is selfdual if $A\leq Z\setminus A$. We will use the notation
@@
We have already introduced Wadge classes in Definition 1.2. Notice that, using the notation of Definition 3.1, the Wadge classes in a space $Z$ are precisely the sets of the form $A\!\!\downarrow\,$ for some $A\subseteq Z$. Also observe that a subset $A$ of a space $Z$ is selfdual iff $A\!\!\downarrow\,$ is selfdual.
@@
The following helpful criterion is a straightforward application of Lemma 3.3. For example, it can be used to show that each $\mathsf{D}_{\eta}(\mathbf{\Sigma}^{0}_{\xi}(Z))$ is a Wadge class whenever $Z$ is an uncountable zero-dimensional Polish space (since all these classes are non-selfdual by [Ke, Exercise 22.26], or by the methods of Chapter 4).
Lemma 3.5.
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be a zero-dimensional Polish space, and let $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(Z)$. If $\mathbf{\Gamma}$ is continuously closed and non-selfdual then $\mathbf{\Gamma}$ is a Wadge class.
The following important result gives the analysis of the selfdual sets (see [CMM2, Corollary 5.5]). In pratice, this result is used to reduce the problem at hand to the non-selfdual case.
Theorem 3.6.
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be a zero-dimensional Polish space, let $A\in\mathbf{\Sigma}(Z)$ be selfdual, and let $U\in\mathbf{\Delta}^{0}_{1}(Z)$. Then there exist pairwise disjoint $U_{n}\in\mathbf{\Delta}^{0}_{1}(U)$ and non-selfdual $A_{n}<A$ in $Z$ for $n\in\omega$ such that $\bigcup_{n\in\omega}U_{n}=U$ and $\bigcup_{n\in\omega}(A_{n}\cap U_{n})=A\cap U$.
Next, we state an elementary result, which shows that clopen sets are “neutral sets” with respect to Wadge-reduction (the straightforward proof is left to the reader). In Chapter 7, we will discuss much more sophisticated closure properties.
@@
We conclude this section with the complete analysis of the non-selfdual Wadge classes below $\mathbf{\Delta}^{0}_{2}$ (see [CMM2, Theorem 11.2]).
@@ 3.2. Relativization
When one tries to give a systematic exposition of Wadge theory, it soon becomes apparent that it would be very useful to be able to say when $A$ and $B$ belong to “the same” Wadge class, even when $A\subseteq Z$ and $B\subseteq W$ for distinct ambient spaces $Z$ and $W$. In the non-selfdual case, this problem can be solved by using Wadge classes in $\omega^{\omega}$ to parametrize Wadge classes in arbitrary zero-dimensional Polish spaces. This is essentially due to Louveau and Saint Raymond (see [LSR2, Theorem 4.2]), but we will follow the more systematic treatment given in [CMM2]. In particular, the following definition and three results establish the foundations of this method (see [CMM2, Section 6] for the proofs).
@@
Until now, we have never assumed the uncountability of the ambient spaces. As the following three results show, the situation gets particularly pleasant when this assumption is satisfied (see [CMM2, Section 7 and Lemma 20.2] for the proofs).
@@
Lemma 3.15.
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be a zero-dimensional Borel space, let $W$ be an uncountable zero-dimensional Polish space, and let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$. Assume that $A\in\mathbf{\Gamma}(Z)$ and $B\!\!\downarrow\,=\mathbf{\Gamma}(W)$. Then there exists a continuous $f:Z\longrightarrow W$ such that $A=f^{-1}[B]$.
@@ 3.3. Level
We begin with some preliminaries concerning the following useful notion.222 According to Wadge, this notion had already been identified by “the classical descriptive set theorists” (see [Wa2, page 187]). Partitioned unions will also play a major role in Chapter 6.
@@
At this point, it is not clear whether every non-selfdual Wadge class $\mathbf{\Gamma}$ even has a level.333 For example, it is conceivable that $\mathsf{PU}_{n}(\mathbf{\Gamma})=\mathbf{\Gamma}$ for all $n<\omega$, while $\mathsf{PU}_{\omega}(\mathbf{\Gamma})\neq\mathbf{\Gamma}$. That this is actually the case is the content of the next theorem (see [CMM2, Corollary 17.2]), whose proof heavily relies on methods from [Lo2].
@@
We conclude this section with a technical result concerning classes of level 00. Lemma 3.22 first appeared as [CMM1, Lemma 14.3], and it will be useful in the proof of Theorem 9.7. By setting $\mathbf{\Gamma}=\mathbf{\Sigma}^{0}_{1}(Z)$ and $A=Z\setminus\{z\}$, where $z\in Z$ is non-isolated, one sees that the assumption that $A$ is codense in $Z$ cannot be dropped.
Lemma 3.22 (Carroy, Medini, Müller).
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $Z$ be an uncountable zero-dimensional Polish space, let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$, and let $A\in\mathbf{\Gamma}$. Assume that $\ell(\mathbf{\Gamma})=0$ and that $A$ is codense in $Z$. Then there exists a non-empty $U\in\mathbf{\Delta}^{0}_{1}(Z)$ and $\mathbf{\Lambda}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$ such that $\mathbf{\Lambda}\subsetneq\mathbf{\Gamma}$ and $A\cap U\in\mathbf{\Lambda}$.
Proof.
Using the assumption that $\ell(\mathbf{\Gamma})=0$, Lemma 3.3 and Proposition 3.17, it is a simple exercise to show that $\widecheck{\mathbf{\Gamma}}\subseteq\mathsf{PU}_{1}(\mathbf{\Gamma})$. This means that we can fix $A_{n}\in\mathbf{\Gamma}$ and $U_{n}\in\mathbf{\Delta}^{0}_{2}(Z)$ for $n\in\omega$ such that $\bigcup_{n\in\omega}U_{n}=Z$, the $U_{n}$ are pairwise disjoint, and $Z\setminus A=\bigcup_{n\in\omega}(A_{n}\cap U_{n})$.
By the Baire category theorem, one can find a non-empty $U\in\mathbf{\Delta}^{0}_{1}(Z)$ such that $U\subseteq U_{n}$ for some $n$. Next, using Lemma 3.7, one sees that $U\setminus A=A_{n}\cap U\in\mathbf{\Gamma}$, hence $Z\setminus(A\cap U)=(U\setminus A)\cup(Z\setminus U)\in\mathbf{\Gamma}$. Therefore, we have $A\cap U\in\mathbf{\Gamma}$ and $Z\setminus(A\cap U)\in\mathbf{\Gamma}$, which easily implies that $\mathbf{\Lambda}=(A\cap U)\!\!\downarrow\,$ has the desired properties if $A\cap U$ is non-selfdual.
To conclude the proof, assume that $A\cap U$ is selfdual. By Theorem 3.6, we can fix non-selfdual $B_{n}<A$ and $V_{n}\in\mathbf{\Delta}^{0}_{1}(Z)$ for $n\in\omega$ such that $\bigcup_{n\in\omega}V_{n}=U$, the $V_{n}$ are pairwise disjoint, and $A\cap U=\bigcup_{n\in\omega}(B_{n}\cap V_{n})$. Since $A$ is codense, there must be $n\in\omega$ such that $V_{n}\neq\varnothing$ and $B_{n}\neq Z$. In particular, it is possible to apply Lemma 3.7, which yields $A\cap V_{n}=B_{n}\cap V_{n}\leq B_{n}$. Setting $\mathbf{\Lambda}=B_{n}\!\!\downarrow\,$ then gives the desired result. ∎
@@ 3.4. Expansions
The following notion can be traced back to [Wa1, Chapter IV], and it is inspired by work of Kuratowski (see [Wa2, Sections 20 and 21]). Once again, our approach will be closer to that of [Lo2], as developed in [CMM2].
@@
[…]
@@
The left-to-right implication is clear by the definition of expansion. In order to prove the other implication, assume that $\mathbf{\Gamma}\nsubseteq\mathbf{\Lambda}$. Then $\widecheck{\mathbf{\Lambda}}\subseteq\mathbf{\Gamma}$ by Lemma  3.3, which implies $\widecheck{\mathbf{\Lambda}}^{(\xi)}\subseteq\mathbf{\Gamma}^{(\xi)}$. So, if we had $\mathbf{\Gamma}^{(\xi)}\subseteq\mathbf{\Lambda}^{(\xi)}$, it would follow that $\mathbf{\Lambda}^{(\xi)}$ is self-dual, contradicting Theorem 3.25. Therefore $\mathbf{\Gamma}^{(\xi)}\nsubseteq\mathbf{\Lambda}^{(\xi)}$, as desired. ∎
@@
whenever $\mathbf{\Gamma},\mathbf{\Gamma}^{(\eta)}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$.
@@
We conclude this section with three technical results. Lemmas 3.29 and 3.30 show that expansions interact in the expected way with the machinery of relativization (see [CMM2, Lemmas 14.1 and 14.2] respectively). Lemma 3.31 is a variation on the theme of Kuratowski’s Transfer Theorem (see [CMM2, Section 12]), and it can be proved like [CMM2, Lemma 14.3].
@@
3.5. Separated differences
Definition 3.32 (Louveau).
Let $Z$ be a space, let $1\leq\eta<\omega_{1}$, and let $U_{\mu,n},A_{\mu,n},A^{\ast}\subseteq Z$ for $\mu<\eta$ and $n\in\omega$. Define
(3.1) $\mathsf{SD}_{\eta}((U_{\mu,n}:\mu<\eta,n\in\omega),(A_{\mu,n}:\mu<\eta,n\in\omega),A^{\ast})=\\ =\bigcup_{\begin{subarray}{c}\mu<\eta\\ n\in\omega\end{subarray}}\left(A_{\mu,n}\cap U_{\mu,n}\setminus\bigcup_{\begin{subarray}{c}\mu^{\prime}<\mu\\ m\in\omega\end{subarray}}U_{\mu^{\prime},m}\right)\cup\left(A^{\ast}\setminus\bigcup_{\begin{subarray}{c}\mu<\eta\\ n\in\omega\end{subarray}}U_{\mu,n}\right).$
Lemma 3.33.
Lemma 3.34.
Separated differences are important because they make it possible to give concrete descriptions of the non-selfdual Wadge classes of level 00. This is made precise by the following result, which is essentially due to Louveau (it can be easily derived using [CMM2, Proposition 19.4 and Theorem 22.2]). See also the first paragraph of Section 6.2 to clarify the meaning of $\mathbf{\Delta}$.
Theorem 3.35.
  • $\mathbf{\Gamma}\in\mathsf{NSD}(Z)$ and $\ell(\mathbf{\Gamma})=0$,
Chapter 4 Wadge theory: Hausdorff operations and universal sets
In this chapter, we will discuss the classical notions of Hausdorff operation and universal set. Ultimately, these notions will allow us to construct new non-selfdual Wadge classes by combinining given ones in a suitable way (see Corollary 4.16). This will be applied to separated unions in Theorem 6.6. Furthermore, the results on universal sets will find a direct application in Lemma 5.4. Along the way, we will discuss the auxiliary notion of relativizable pointclass. Finally, we will show how to interpret level and expansions in the context of Hausdorff operations.
4.1. Hausdorff operations
While the following notion is certainly not the focus of this article, it is technically very useful, since it captures the standard way to produce non-selfdual Wadge classes. In fact, in the next section, we will show that classes with a universal set must be non-selfdual, and Hausdorff operations are an appropriate tool to manipulate universal sets. The exposition in this section is based on [CMM2, Sections 8 and 9], and all results are straightforward to prove except Theorem 4.7.
@@ 4.1. Hausdorff operations
Notice that the set $D$ in the above definition simultaneously yields a Hausdorff operation for every set $Z$. In case it will be important to specify the ambient set, we will use the notation $\mathcal{H}_{D}^{Z}$.
A basic fact about Hausdorff operations, expressed by the following proposition, is that they include all operations obtained by combinining unions, intersections and complements.
Proposition 4.2.
Let $I$ be a non-empty set, and let $D_{i}\subseteq\mathcal{P}(\omega)$ for $i\in I$. Fix an ambient set $Z$ and $A_{0},A_{1},\ldots\subseteq Z$.
  • $\mathcal{H}_{S_{n}}(A_{0},A_{1},\ldots)=A_{n}$ for all $n\in\omega$, where $S_{n}=\{A\subseteq\omega:n\in A\}$.
  • $\bigcap_{i\in I}\mathcal{H}_{D_{i}}(A_{0},A_{1},\ldots)=\mathcal{H}_{D}(A_{0},A_{1},\ldots)$, where $D=\bigcap_{i\in I}D_{i}$.
  • $\bigcup_{i\in I}\mathcal{H}_{D_{i}}(A_{0},A_{1},\ldots)=\mathcal{H}_{D}(A_{0},A_{1},\ldots)$, where $D=\bigcup_{i\in I}D_{i}$.
  • $Z\setminus\mathcal{H}_{D}(A_{0},A_{1},\ldots)=\mathcal{H}_{\mathcal{P}(\omega)\setminus D}(A_{0},A_{1},\ldots)$ for all $D\subseteq\mathcal{P}(\omega)$.
Furthermore, the composition of Hausdorff operations still yields a Hausdorff operation, as described by the following proposition.
Proposition 4.3.
Let $Z$ be a set, let $D\subseteq\mathcal{P}(\omega)$ and $E_{n}\subseteq\mathcal{P}(\omega)$ for $n\in\omega$. Fix a bijection $\sigma:\omega\times\omega\longrightarrow\omega$. Then there exists $F\subseteq\mathcal{P}(\omega)$ such that
$\mathcal{H}_{D}(B_{0},B_{1},\ldots)=\mathcal{H}_{F}(A_{0},A_{1},\ldots)$
for all $A_{0},A_{1},\ldots\subseteq Z$, where $B_{n}=\mathcal{H}_{E_{n}}(A_{\sigma(n,0)},A_{\sigma(n,1)},\ldots)$ for $n\in\omega$.
As an application, one can combine Propositions 4.2 and 4.3 to prove by induction that each $\mathbf{\Sigma}^{0}_{\xi}(Z)$ is a Hausdorff class for every space $Z$. Furthermore, the same will be true for the classes $\mathsf{D}_{\eta}(\mathbf{\Sigma}^{0}_{\xi}(Z))$.
The following result shows that Hausdorff operations are well-behaved with respect to inverse images and subsets.
Lemma 4.4.
Let $D\subseteq\mathcal{P}(\omega)$, let $Z$ and $W$ be sets, let $A_{0},A_{1},\ldots\subseteq Z$, and let $B_{0},B_{1},\ldots\subseteq W$.
  • $f^{-1}[\mathcal{H}_{D}(B_{0},B_{1},\ldots)]=\mathcal{H}_{D}(f^{-1}[B_{0}],f^{-1}[B_{1}],\ldots)$ for all $f:Z\longrightarrow W$.
  • $W\cap\mathcal{H}_{D}^{Z}(A_{0},A_{1},\ldots)=\mathcal{H}_{D}^{W}(A_{0}\cap W,A_{1}\cap W,\ldots)$ whenever $W\subseteq Z$.
In the present context, the significance of the following definition lies in the fact that, under suitable determinacy assumptions, it gives an alternative description of non-selfdual Wadges classes (see Theorem 4.7).
@@
The following result is the analogue of relativization for Hausdorff classes, and it can be proved using Lemma 4.4.
Lemma 4.6.
Let $D\subseteq\mathcal{P}(\omega)$, and let $Z$ and $W$ be spaces.
  • If $f:Z\longrightarrow W$ is continuous and $B\in\mathbf{\Gamma}_{D}(W)$ then $f^{-1}[B]\in\mathbf{\Gamma}_{D}(Z)$.
  • Assume that $W\subseteq Z$. Then $\mathbf{\Gamma}_{D}(W)=\{A\cap W:A\in\mathbf{\Gamma}_{D}(Z)\}$.
For a proof of the following result, see [CMM2, Theorem 22.2]. The case $Z=\omega^{\omega}$ and $\mathbf{\Sigma}=\mathcal{P}$ is originally due to Van Wesep (see [VW1, Theorem 5.3.1]).
@@
  • $\mathbf{\Gamma}=\mathbf{\Gamma}_{D}(Z)$ for some $D\subseteq\mathcal{P}(\omega)$.
4.2. Relativizable pointclasses and universal sets
In Section 2.2, we have seen that nice topological pointclasses provide a suitable framework to formulate determinacy assumptions. Next, we will give another application of topological pointclasses, by showing that the following notion provides a suitable framework to formulate and prove results concerning universal sets.
Definition 4.8.
  • $\mathbf{\Pi}$ is a topological pointclass,
Given $D\subseteq\mathcal{P}(\omega)$, using Lemma 4.6, one sees that setting $\mathbf{\Pi}(Z)=\mathbf{\Gamma}_{D}(Z)$ for every space $Z$ yields a relativizable pointclass. However, in the present context, the most important example is given by the following definition. Notice that the assignment described by Definition 4.9 (under suitable determinacy assumptions) gives a relativizable pointclass by Lemmas 3.11 and 3.12. In fact, the reason for the restriction “zero-dimensional Borel” in Definitions 4.8 and 4.9 is that Lemma 3.12 only applies to such spaces.
Definition 4.9.
Let $\mathbf{\Gamma}\in\mathsf{NSD}(\omega^{\omega})$. We will say that a relativizable pointclass $\mathbf{\Pi}$ induced by $\mathbf{\Gamma}$ if $\mathbf{\Pi}(Z)=\mathbf{\Gamma}(Z)$ for every zero-dimensional Borel space $Z$.
The following is a classical notion (see for example [Ke, Definition 22.2]).
@@ 4.2. Relativizable pointclasses and universal sets
Next, we will prove three results that are essentially the same as [CMM2, Proposition 10.2, Corollary 10.3, and Lemma 10.4], except that we will formulate them in the more general framework described above. We remark that Proposition 4.11 could be viewed as a warm-up for Theorem 4.15.
@@
By [Ke, Theorem 22.3], we can fix a $2^{\omega}$-universal set $U$ for $\mathbf{\Sigma}^{0}_{1}(Z)$. Let $h:2^{\omega}\longrightarrow(2^{\omega})^{\omega}$ be a homeomorphism, and let $\pi_{n}:(2^{\omega})^{\omega}\longrightarrow 2^{\omega}$ denote the projection on the $n$-th coordinate for $n\in\omega$. Set $(y)_{n}=\pi_{n}(h(y))$ for $y\in 2^{\omega}$ and $n\in\omega$. Pick a space $Z$. By [Ke, Theorem 22.3], we can fix a $2^{\omega}$-universal set $U$ for $\mathbf{\Sigma}^{0}_{1}(Z)$. Define
$U_{n}=\{(y,x)\in 2^{\omega}\times Z:((y)_{n},x)\in U\}$
for $n\in\omega$, and observe that each $U_{n}\in\mathbf{\Sigma}^{0}_{1}(2^{\omega}\times Z)$.
Set $S=\mathcal{H}_{D}(U_{0},U_{1},\ldots)$. In order to conclude the proof, we will show that $S$ is a $2^{\omega}$-universal set for $\mathbf{\Gamma}_{D}(Z)$. Using Lemma 4.6, it is easy to see that $S\in\mathbf{\Gamma}_{D}(2^{\omega}\times Z)$ and $S_{y}\in\mathbf{\Gamma}_{D}(Z)$ for every $y\in 2^{\omega}$. Pick $A\in\mathbf{\Gamma}_{D}(Z)$. Let $A_{n}\in\mathbf{\Sigma}^{0}_{1}(Z)$ for $n\in\omega$ be such that $A=\mathcal{H}_{D}(A_{0},A_{1},\ldots)$. Since $U$ is $2^{\omega}$-universal, we can fix $y_{n}\in 2^{\omega}$ for $n\in\omega$ such that $U_{y_{n}}=A_{n}$ for each $n$. Let $y\in 2^{\omega}$ be defined by requiring that $(y)_{n}=y_{n}$ for each $n$. It is straightforward to verify that $S_{z}=A$. ∎
@@
Let $\mathbf{\Pi}$ be a relativizable topological pointclass, and let $Z$ be an uncountable zero-dimensional Borel space. If there exists a $2^{\omega}$-universal set for $\mathbf{\Pi}(Z)$ then there exists a $Z$-universal set for $\mathbf{\Pi}(Z)$.

### v3

@@ 3.1. The basics of Wadge theory
We begin with the most fundamental definitions and results of Wadge theory. For the proofs, see [CMM2, §4]. Assume that a set $Z$ and $\mathbf{\Gamma}\subseteq\mathcal{P}(Z)$ are given. We will use the notation $\widecheck{\mathbf{\Gamma}}=\{Z\setminus A:A\in\mathbf{\Gamma}\}$ for the dual of $\mathbf{\Gamma}$, and say that $\mathbf{\Gamma}$ is selfdual if $\mathbf{\Gamma}=\widecheck{\mathbf{\Gamma}}$. Also set $\Delta(\mathbf{\Gamma})=\mathbf{\Gamma}\cap\widecheck{\mathbf{\Gamma}}$.
@@
Let $Z$ be a space, and let $A,B\subseteq Z$. We will write $A\leq B$ if there exists a continuous function $f:Z\longrightarrow Z$ such that $A=f^{-1}[B]$. In this case, we will say that $A$ is Wadge-reducible to $B$, and that $f$ witnesses the reduction.999 Wadge-reduction is usually denoted by $\leq_{\mathsf{W}}$, which allows to distinguish it from other types of reduction (such as Lipschitz-reduction). Since we will not consider any other type of reduction in this article, we decided to simplify the notation. We will write $A<B$ if $A\leq B$ and $B\not\leq A$. We will say that $A$ is selfdual if $A\leq Z\setminus A$. We will use the notation
@@
We have already introduced Wadge classes in Definition 1.2. Notice that, using the notation of Definition 3.1, the Wadge classes in a space $Z$ are precisely the sets of the form $A\!\!\downarrow\,$ for some $A\subseteq Z$.
@@
Next, we state an elementary result, which shows that clopen sets are “neutral sets” with respect to Wadge-reduction (the straightforward proof is left to the reader). In §5, we will discuss much more sophisticated closure properties.
@@
We conclude this subsection with the complete analysis of the non-selfdual Wadge classes below $\mathbf{\Delta}^{0}_{2}$ (see [CMM2, Theorem 11.2]).
@@ 3.2. Relativization
When one tries to give a systematic exposition of Wadge theory, it soon becomes apparent that it would be very useful to be able to say when $A$ and $B$ belong to “the same” Wadge class, even when $A\subseteq Z$ and $B\subseteq W$ for distinct ambient spaces $Z$ and $W$. In the non-selfdual case, this problem can be solved by using Wadge classes in $\omega^{\omega}$ to parametrize Wadge classes in arbitrary zero-dimensional Polish spaces. This is essentially due to Louveau and Saint Raymond (see [LSR2, Theorem 4.2]), but we will follow the more systematic treatment given in [CMM2]. In particular, the following definition and three results establish the foundations of this method (see [CMM2, §6] for the proofs).
@@
Until now, we have never assumed the uncountability of the ambient spaces. As the following two results show, the situation gets particularly pleasant when this assumption is satisfied (see [CMM2, §7 and Lemma 20.2] for the proofs).
@@ 3.3. Level
We begin with some preliminaries concerning the following useful notion.101010 According to Wadge, this notion had already been identified by “the classical descriptive set theorists” (see [Wa2, page 187]). Partitioned unions will also play a major role in §4.
@@
At this point, it is not clear whether every non-selfdual Wadge class $\mathbf{\Gamma}$ even has a level.111111 For example, it is conceivable that $\mathsf{PU}_{n}(\mathbf{\Gamma})=\mathbf{\Gamma}$ for all $n<\omega$, while $\mathsf{PU}_{\omega}(\mathbf{\Gamma})\neq\mathbf{\Gamma}$. That this is actually the case is the content of the next theorem (see [CMM2, Corollary 17.2]), whose proof heavily relies on methods from [Lo2].
@@ 3.4. Expansions
The following notion can be traced back to [Wa1, Chapter IV], and it is inspired by work of Kuratowski (see [Wa2, §20 and §21]). Once again, our approach will be closer to that of [Lo2], as developed in [CMM2].
@@
$\mathbf{\Gamma}^{(\xi)}=\{f^{-1}[A]:A\in\mathbf{\Gamma}\text{ and }f:Z\longrightarrow Z\text{ is $\mathbf{\Sigma}^{0}_{1+\xi}$-measurable}\}.$
@@
The left-to-right implication is clear by the definition of expansion. In order to prove the other implication, assume that $\mathbf{\Gamma}\nsubseteq\mathbf{\Lambda}$. Then $\widecheck{\mathbf{\Lambda}}\subseteq\mathbf{\Gamma}$ by Lemma 3.3, which implies $\widecheck{\mathbf{\Lambda}}^{(\xi)}\subseteq\mathbf{\Gamma}^{(\xi)}$. So, if we had $\mathbf{\Gamma}^{(\xi)}\subseteq\mathbf{\Lambda}^{(\xi)}$, it would follow that $\mathbf{\Lambda}^{(\xi)}$ is self-dual, contradicting Theorem 3.21. Therefore $\mathbf{\Gamma}^{(\xi)}\nsubseteq\mathbf{\Lambda}^{(\xi)}$, as desired. ∎
@@
whenever $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(Z)$.
@@
We conclude this subsection with three technical results. Lemmas 3.25 and 3.26 show that expansions interact in the expected way with the machinery of relativization (see [CMM2, Lemmas 14.1 and 14.2] respectively). Lemma 3.27 is a variation on the theme of Kuratowski’s Transfer Theorem (see [CMM2, §12]), and it can be proved like [CMM2, Lemma 14.3].
@@
3.5. Hausdorff operations and universal sets
The main purpose of this subsection is to obtain Theorems 3.34 and 3.37, which explain the relationship between universal sets and non-selfdual Wadge classes. This results will be useful in the proofs of Corollary 3.39 and Lemma 3.48. One of the fundamental tools to prove results of this kind is given by the notion of Hausdorff operation.
@@ 3.5. Hausdorff operations and universal sets
A basic fact about Hausdorff operations is that they include all operations obtained by combinining unions, intersections and complements (see [CMM2, Proposition 8.2]). Furthermore, the composition of Hausdorff operations is still a Hausdorff operation (see [CMM2, Proposition 8.3]). As an application, one can use induction to prove that each $\mathbf{\Sigma}^{0}_{\xi}(Z)$ is a Hausdorff class. In fact, as the following result shows, something much more general holds (see [CMM2, Theorem 22.2]). The case $Z=\omega^{\omega}$ and $\mathbf{\Sigma}=\mathcal{P}$ is originally due to Van Wesep (see [VW1, Theorem 5.3.1]).
@@
  • $\mathbf{\Gamma}=\mathbf{\Gamma}_{D}(Z)$ for some $D\subseteq\mathcal{P}(\omega)$ and $\mathbf{\Gamma}\subseteq\mathbf{\Sigma}(Z)$.
Next, we come to the classical notion of a universal set (see for example [Ke, Definition 22.2]).
@@
The following special kind of topological pointclass is particularly useful when dealing with universal sets.
Definition 3.32.
  • $\mathbf{\Pi}$ is a topological pointclass,
In the present context, the relevant examples of relativizable pointclasses are obtained as follows:
  • Given $D\subseteq\mathcal{P}(\omega)$, set $\mathbf{\Pi}(Z)=\mathbf{\Gamma}_{D}(Z)$ for every space $Z$,
  • Given $\mathbf{\Gamma}\in\mathsf{NSD}(\omega^{\omega})$, set $\mathbf{\Pi}(Z)=\mathbf{\Gamma}(Z)$ for every space $Z$.
In the first case, this is the content of [CMM2, Lemma 9.5]. In the second case (under suitable determinacy assumptions), this follows from Lemmas 3.9 and 3.10. In fact, the reason for the restriction “zero-dimensional Borel” in Definition 3.32 is that Lemma 3.10 only applies to such spaces.
Next, we will see that non-selfdual Wadge classes have universal sets. We will need the following auxiliary result (see [CMM2, Proposition 10.2] for the proof).
@@
Theorem 3.34.
Let $\mathbf{\Sigma}$ be a nice topological pointclass, and assume that $\mathsf{Det}(\mathbf{\Sigma}(\omega^{\omega}))$ holds. Let $\mathbf{\Gamma}\in\mathsf{NSD}_{\mathbf{\Sigma}}(\omega^{\omega})$, and let $Z$ be a zero-dimensional Borel space. Then $\mathbf{\Gamma}(Z)$ has a $2^{\omega}$-universal set.
@@
By Theorem 3.30, we can fix $D\subseteq\mathcal{P}(\omega)$ such that $\mathbf{\Gamma}=\mathbf{\Gamma}_{D}(\omega^{\omega})$. Observe that if $W$ is a zero-dimensional Borel space, which we assume without loss of generality to be a subspace of $\omega^{\omega}$, then
$\mathbf{\Gamma}(W)=\{A\cap W:A\in\mathbf{\Gamma}(\omega^{\omega})\}=\{A\cap W:A\in\mathbf{\Gamma}_{D}(\omega^{\omega})\}=\mathbf{\Gamma}_{D}(W)$
by relativizability. In other words, the same $D$ works for every zero-dimensional Borel space. To conclude the proof, simply apply Proposition 3.33. ∎
Finally, we will obtain a converse to Theorem 3.34. We will need the following two auxiliary results, which correspond to [CMM2, Corollary 10.3 and Lemma 10.4], except that they are formulated in the more general framework described above.
@@
Let $\mathbf{\Pi}$ be a relativizable topological pointclass, and let $Z$ be an uncountable zero-dimensional Borel space. If $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set then $\mathbf{\Pi}(Z)$ has a $Z$-universal set.
Lemma 3.36.
Let $\mathbf{\Pi}$ be a topological pointclass, and let $Z$ be a space. If $\mathbf{\Pi}(Z)$ has a $Z$-universal set then $\mathbf{\Pi}(Z)$ is non-selfdual.
Theorem 3.37.
Let $\mathbf{\Pi}$ be a relativizable pointclass, and let $Z$ be an uncountable zero-dimensional Borel space. If $\mathbf{\Pi}(Z)$ has a $2^{\omega}$-universal set then $\mathbf{\Pi}(Z)\in\mathsf{NSD}(Z)$.